65) Annika Company uses activity-based costing. The company has two products: A and B. The

annual production and sales of Product A is 4,000 units and of Product B is 1,000 units. There

are three activity cost pools, with total cost and activity as follows:

Total Activity

Activity Cost Pool Total Cost Product A Product B Total

Activity 1 $ 18,000 700 300 1,000

Activity 2 $ 24,000 500 100 600

Activity 3 $ 60,000 800 400 1,200

The activity-based costing cost per unit of Product A is closest to:

A) $20.40

B) $10.00

C) $18.15

D) $17.00


1
Expert's answer
2021-04-26T19:21:23-0400

Solution:

The activity-based costing cost per unit of Product A is closest to C). 18.15

 

Activity cost of Product A = "\frac{Total Activity}{Total Cost}"


Activity 1 = "\frac{700}{1000}\times 18,000 = \$12,600"


Activity 2 = "\frac{500}{600}\times 24,000 = \$20,000"


Activity 3 ="\frac{800}{1200}\times 60,000 = \$40,000"


Total cost of Product A = 12,600 + 20,000 + 40,000 = $72,600


To derive the cost per unit of Product A, we divide the total cost by the annual production and sales of Product A, which is 4,000 units.


= "\frac{72,600}{4000} = \$18.15"


The cost per unit of Product A = $18.15